OpenStudy (anonymous):

So I need to arrange 4 science books and 3 math books in such manner that science books are together and math books are together laying on top va each other. How many ways I can do this?

OpenStudy (anonymous):

Me thinks 4*3*2*1+3*2*1=30

OpenStudy (anonymous):

Any thoughts?

OpenStudy (anonymous):

@DaBest21

OpenStudy (misssunshinexxoxo):

This is a combination. What is 4 times 3?

OpenStudy (anonymous):

12

OpenStudy (misssunshinexxoxo):

That is the number of different ways you can change the way they are situated

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

boo

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

crying face

OpenStudy (solomonzelman):

This is what I think. There are 4P4 ways to rearrange the science books, AND there are 3P3 ways to rearrange the math books. And all books have to be on top of each other, would mean either math books on top of science books, or science books on top of math books. that means that all your previous permutations should be multiplied times two.

OpenStudy (solomonzelman):

it should end up pretty big (not 12 or 11)

OpenStudy (anonymous):

4*3*2*1=24 3*2*1=6

OpenStudy (anonymous):

so 30

OpenStudy (anonymous):

or 60

OpenStudy (anonymous):

Ok but answer says 288!!!!!

OpenStudy (anonymous):

no entiendo por nada!

OpenStudy (solomonzelman):

yes, because you multiply the permuations by math books, times the permutations by science books

OpenStudy (solomonzelman):

and then multiply by two, because math books can be on top of sceince books or vv

OpenStudy (solomonzelman):

\(\left({\rm 4P4}\times {\rm 3P3}\right)\times 2\)

OpenStudy (solomonzelman):

for each permutation of math books, there is all permutations of science books, and vice versa. that is why you multiply.

OpenStudy (solomonzelman):

and then ×2, because either math books can be on top of science, or the other way around.

OpenStudy (anonymous):

You rock!

OpenStudy (anonymous):

I need to be careful who I ask

OpenStudy (anonymous):

people misinformed all the time

OpenStudy (solomonzelman):

Lol, not really....:) thanks tho:)

OpenStudy (solomonzelman):

Yeah, sometimes answer is wrong, but you have to do a little analysis if the answer makes sense or not.

OpenStudy (solomonzelman):

In this case, for example. How can answer be 11 or 12, if the number of combinations of just the science books alone exceeds that?

OpenStudy (anonymous):

Yeah. I was nearly running away with 12 which would be a not so correct but intuitive answer.

OpenStudy (solomonzelman):

see what I posted about 11 and 12 just now?

OpenStudy (anonymous):

Yeah. Permutations

OpenStudy (solomonzelman):

i will rephrase, there is a typo....

OpenStudy (solomonzelman):

combinations should be "permutations"

OpenStudy (solomonzelman):

But, in any case, now you know.... check your answer to make sense:)

OpenStudy (solomonzelman):

have a good one!

OpenStudy (anonymous):

Thanks so much! I'll be writing about you